OpenStudy (anonymous):

please help me..... how do i simplify (sqrt(3)-i)^9

OpenStudy (anonymous):

\[(\sqrt{3}-i)^{9}= [(\sqrt{3}-i)^{3}]^{3}\] \[(\sqrt{3}-i)^{3}= (\sqrt{3}-i)(\sqrt{3}-i)(\sqrt{3}-i)\]

OpenStudy (anonymous):

I wonder how to simplify it... :(

OpenStudy (anonymous):

First, u calculate\[(\sqrt{3}-i)^{2}=(\sqrt{3}-i)(\sqrt{3}-1)\] then, u use what u got to calculate \[(\sqrt{3}-i)^{2}(\sqrt{3}-i)= (\sqrt{3}-i)^{3}\] let \[A=(\sqrt{3}-i)^{3}\] \[(\sqrt{3}-i)^{9}= [(\sqrt{3}-i)^{3}]^{3} = A ^{3}\]

OpenStudy (anonymous):

remember that i*i=-1